Find the distance between (–3, –2) and (1, 4)
Answer by checkley75(3666)

FIRST WE NEED TO FIND THE DISTANCE BETWEEN THE X & Y POINTS.
X=(1+3)=4
Y=(4+2)=6
NOW THAT WE HAVE TWO OF THE SIDES WE NEED TO FIND THE DIAGONAL USING THE PATHAGOREAN FORMULA A^2+B^2=C^2
4^2+6^2=C^2
16+36=C^2
52=C^2
C=7.21 ANSWER.